# Center for Nonlinear Studies, Los Alamos National Laboratory, Los Alamos, New Mexico 87545

### v1

@@
Center for Nonlinear Studies, Los Alamos National Laboratory, Los Alamos, New Mexico 87545
@@
Sakib Matin sakibmatin@gmail.com Theoretical Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Emily Shinkle Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Yulia Pimonova Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Galen T. Craven Theoretical Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Aleksandra Pachalieva Earth and Environmental Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Ying Wai Li Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Kipton Barros Theoretical Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Nicholas Lubbers Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA
@@
Machine learning interatomic potentials (MLIPs) are a promising tool to accelerate atomistic simulations and molecular property prediction. The quality of MLIPs strongly depends on the quantity of available training data as well as the quantum chemistry (QC) level of theory used to generate that data. Datasets generated with high-fidelity QC methods, such as coupled cluster, are typically restricted to small molecules and may be missing energy gradients. With this limited quantity of data, it is often difficult to train good MLIP models. We present an ensemble knowledge distillation (EKD) method to improve MLIP accuracy when trained to energy-only datasets. In our EKD approach, first, multiple teacher models are trained to QC energies and then used to generate atomic forces for all configurations in the dataset. Next, a student MLIP is trained to both QC energies and to ensemble-averaged forces generated by the teacher models. We apply this workflow on the ANI-1ccx dataset which consists of organic molecules with configuration energies computed at the coupled cluster level of theory. The resulting student MLIPs achieve new state-of-the-art accuracy on the out-of-sample COMP6 benchmark and improved stability for molecular dynamics simulations. The EKD approach for MLIP is broadly applicable for chemical, biomolecular and materials science simulations.
Machine learning models are a promising way to accelerate scientific simulations 1, 2, 3, 4. Machine learning interatomic potentials (MLIPs) 1, 2, 3 can emulate the potential energy surface of atomistic systems at dramatically reduced costs compared to reference quantum chemistry (QC) algorithms. MLIPs 5, 6, 2, 1, 3, 7 are typically trained on QC datasets to map from atomic coordinates and species to configuration energy. The atomic forces can be predicted by the trained MLIPs, using automatic differentiation. Generating the data used to train MLIPs is far more computationally expensive than performing the training itself 8. It is especially computationally difficult to generate training data using high fidelity QC methods, such as Coupled Cluster, Configuration-Interaction, Quantum Monte Carlo, etc, because computational cost typically explodes with the number of electrons 9. Because of the much higher costs relative to density functional theory (DFT) calculations, datasets of high-fidelity QC are typically small, both in the size and number of molecular configurations. Furthermore, many high-fidelity QC codes provide only the total energy, but no gradients, which further hinders training of MLIPs 10. Our work focuses on techniques to better utilize such small high fidelity QC datasets for training MLIPs.
@@
There has been tremendous progress in the field of MLIP development, especially in terms of designing more expressive architectures, including equivariant descriptors 11, 12, 13 and transformer-based architectures 14. Furthermore recent works have focused on generating larger training data sets 15, and scalable training 16. On the other hand, an under-explored area is the design of better training protocols, especially for low data regimes. Transfer learning 17 and multi-fidelity learning 18 can be effective for small high fidelity datasets but they may require large amounts of data at a different level of theory to be successful.
@@
Recently, Knowledge Distillation (KD) for MLIPs 19, 20 has been shown to be an effective training protocol for existing datasets, without expensive pre-training. In the prototypical KD workflow 21, a single teacher model generates auxiliary outputs that augment the training of a student model in order to enhance speed 22, 19, memory requirements 23, and accuracy 24. References 25, 24, 26 have shown that multiple teachers can train a single student model to improve performance in classification tasks. KD has been applied to MLIPs to accelerate molecular dynamics (MD) simulations, by using intermediate outputs (e.g., atomic energies) 20, learned features 19, and hessians of the energy 27. Similar teacher-student training has also been applied to materials structure for property prediction tasks 28. In related works (see Refs. 29, 30, 31), the teacher MLIP (trained on QC ground truth data) is used to generate synthetic data by running MD under different conditions. The student MLIPs is first trained to the synthetic data 29, 30, and optionally fine-tuned on other QC ground truth data 31. Existing studies have mostly focus on utilizing a single teacher and student workflow, even though using ensemble-averaged MLIPs in MD simulations has been shown to improve accuracy and stability 32. Only, Ref. 33 studied knowledge distillation with an ensemble five graph neural networks (GNNs) trained on energy, forces and stress, which were evaluated at DFT level of theory. A GNN was chosen at random to generate new trajectories, for which the ensemble was used to predict energy, forces and stress, and then the model was fine-tuned to the mean predictions 33.
@@
Distinct from previous publications 19, 28, 33, 20, in our Ensemble Knowledge Distillation (EKD) for MLIPs workflow, a set of teacher MLIPs are trained on high fidelity data that only contains molecular energies. The trained teacher models can generate the atomic forces for all configurations in the dataset. The students are trained to the ground truth QC energies and the ensemble averaged forces from the teachers. Our workflow is outlined in Fig 1. We validate our EKD workflow on the ANI-1ccx dataset 17, 34 to show student models are more accurate and robust than direct training. Our workflow establishes a new state-of-the-art accuracy for the ANI-1ccx dataset without using any expensive pre-training 17.
@@
Figure 1: Ensemble Knowledge Distillation for Hierarchically Interacting Particle Neural Network (HIPNN) 35, 13. The $k$ teacher models are trained on the reference quantum chemistry calculation of the energy and can generate forces (negative of the gradient of energy with respect to position). These forces are averaged over the ensemble of teachers and augment the student HIPNN training to improve the accuracy and robustness.
@@
In this manuscript, we apply the EKD method primarily to the Hierarchically Interacting Particle Neural Network (HIPNN) architecture 35, although this method can be readily applied to most existing MLIPs. HIPNN is a message-passing graph neural network 7, that can map atomic configurations to energy 35, forces 13, and to various chemical properties 36, 37, 38. HIPNN uses one-hot encoding, based on the atomic number, to initially featurize the local atomic environments. Then the interaction layers allow for mixing of atomic environments between neighbors (within a local cutoff) via message passing to refine the initial features 35. Using multiple interaction layers, $n_{\mathrm{Int}}>1$, implicitly accounts for some long-range physics 35, 13, 7. Although HIPNN 35 originally utilized scalar pair-wise distances between neighbors, the subsequent generalization—HIPNN with tensor sensitivity 13—utilizes higher order Cartesian tensor products of the displacement vectors between neighboring atoms to construct more informative many-body descriptors. The hyper-parameter $l_{\mathrm{max}}$ corresponds to the highest order tensor used in the model. For $\ell_{\mathrm{max}}>0$, the model predictions are sensitive to the angles between neighboring atoms. The $\ell_{\mathrm{max}}=0$ HIPNN model coincides with the model developed in original publication 35. Finally, the atom-layers (multi-layer perceptrons) predict the atomic energy $\mathcal{\epsilon}_{i}$ which are then summed up to obtain the configuration, $E$. Automatic differentiation can be used to compute the forces on each atom $\boldsymbol{F}_{i}=\boldsymbol{\nabla}_{i}E$. The hyper-parameters for the HIPNN models are given in Supporting Information.  1.1.
@@
We validate the developed EKD workflow on the ANI-1ccx 17, 34 dataset, which consists of small organic molecules. The approximately $4.9\times 10^{5}$ molecular configurations in this dataset span $\mathrm{C,H,N,O}$ elements. The configurations have been down selected from the larger ANI-1x dataset (about 5 million configurations) using active learning 17, which utilized ensemble disagreement as the uncertainty metric. The non-equilibrium geometries are generated via normal mode sampling and short MD trajectories 39. The ANI-1ccx datasets include additional dihedral sampling of small molecules that is not present in the ANI-1x dataset. The dataset is available for download from Ref. 34. The configuration energies are computed at the coupled cluster with singles doubles, perturbative-triples, and complete basis set extrapolation [CCSD(T)/CBS] level of theory 40, 17 using the ORCA software 41. The dataset has also been computed at the DFT level of theory using the $\omega\text{B}97x$ functional, which is the same level of theory as the ANI-1x dataset 39.
@@
The COMP6 39, 17 is a challenging out-of-sample test for models trained to the ANI-1ccx and ANI-1x datasets. The configurations in the COMP6 are larger than the ANI-1x and ANI-1ccx training datasets, and provide a challenging extensibility test. For the ‘torsion’ 42 and ‘GDB 10-13’ data subsets, the energy and conformer energy differences have been computed at the CCSD(T)/CBS level of theory and we refer to these as the CC-COMP6 dataset. The conformational energy $\Delta E$ is the energy difference between all conformers for a given molecule in the benchmark. The conformers with energies at least $100\mathrm{kcal/mol}$ greater than the ground state are excluded, similar to the analysis in Ref. 17.
@@
Figure 2: Student HIPNNs have lower root-mean-squared-errors (RMSE) for energy $E$ and conformer energy differences $\Delta E$ compared to the teacher models. The out-of-sample test sets “GDB 10-13” and “Torsion” 42 are subsets of the CC-COMP6 dataset 34.
@@
which contains the atomic positions $\boldsymbol{R}_{i}$ and species $Z_{i}$ for each configuration and the corresponding energy $E$. All $8$ teacher models have the same architecture, but are initialized with different random weights and different data splits. We train the HIPNN models using stochastic gradient descent. The loss function consists of both error and regularization terms. The error loss $\mathcal{L}_{\mathrm{err}}$ is the sum of the root-mean-squared error (RMSE) and mean-absolute error (MAE) losses. The regularization term consists of the $L_{2}$ norm of model weights $\mathcal{L}_{L_{2}}$, which is commonly added to loss functions to reduce over-fitting, and the hierarchicality term $\mathcal{L}_{\mathrm{R}}$ term 35, which is specific to HIPNN.
@@
Although the teacher models are trained only on molecular energies, they can predict the forces using automatic differentiation. The ensemble-averaged teacher forces for each atom $i$,
@@
We use a loss scheduler where the $w_{\mathrm{F}}$ is dynamically updated during training. The value of $w_{\mathrm{F}}$ is larger during the early stages and slowly decreases to a smaller value. The student model training benefits from the local information provided by the $\overline{\boldsymbol{F}}_{i}$ in the early stages of training, followed by refinement on the QC configuration energies at the final stages. Similar loss scheduler 43 strategies have shown improved accuracy when training MLIPs. The weights of the loss function and scheduler are listed in the Supplementary Information 1.2. The force predictions from an ensemble of the first generation student models can be used to train the second generation of student HIPNNs.
We apply the EKD workflow to HIPNNs trained on the ANI-1ccx data set. The student HIPNNs achieve lower root-mean-squared error (RMSE) for energy $E$ and conformer energy differences $\Delta E$ in the out-of-sample CC-COMP6 benchmark compared to the set of teacher models in Fig. 2. The average error (standard deviation) of the student HIPNNs is lower than the error of the ensembled predictions of teacher models. The ensembled teachers are more accurate than single teachers, but they have slower MD speed and increased memory requirements. By using EKD, we get a single student that does just as well or even better than the ensembled teacher. Thus, we can get all the gains in accuracy from the ensembled teacher model without the computational disadvantages of ensembled models.
Figure 3: MD simulations driven by student HIPNNs are more stable than teacher models. The student and teachers are the same models as in Fig. 2. The MD simulations are performed with individual models, not ensembled predictions.
@@
To investigate how our EKD workflow is affected by the capacity gap between the teacher and student MLIPs, we vary the student models’ number of interaction layers $n_{\mathrm{Int}}$ and tensor sensitivity order $\ell_{\mathrm{max}}$. Increasing $n_{\mathrm{Int}}$ and $l_{\mathrm{max}}$ means that descriptors are more sensitive to many-body angular information about neighboring atoms, i.e., the resulting models have greater capacity 13. We fix the the width of the neural network, depth of the atom layers, and number of sensitivity functions, as well as the training hyper-parameters (number of epochs, learning rates, and optimizer) for all models considered. The control models have the same $n_{\mathrm{Int}}$ and $\ell_{\mathrm{max}}$ as the students but they are trained only on the QC energies. We set $n_{\mathrm{Int}}=2,\ell_{\mathrm{max}}=2$ for the teacher HIPNNs. Figure 4 shows that $E$ and $\Delta E$ RMSE of the student HIPNNs are consistently 10-30% lower than the control models on the CC-COMP6 benchmark.
We implement the EKD method with different MLIP architectures for teachers and students. We use TorchANI-1ccx MLIPs from Refs. 45 and 46, which were first pre-trained to the nearly 5 million configurations of the DFT ANI-1x dataset, and fine-tuned to the coupled cluster ANI-1ccx dataset 17. We compare the $\Delta E$ RMSE of the “GDB 10-13” subset of the CC-COMP6 benchmark in Fig. 5. We find that the using a teacher torchANI model does benefit a student HIPNN model, compared to direct training (teacher HIPNNs). Ultimately, we find that using HIPNN MLIPs for both teacher and student gives lowest error, which we attribute to the fact the teacher HIPNNs have lower errors than the teacher TorchANI models.
@@
The accuracy of the forces of the student models in our EKD workflow is an important metric to analyze because of the strong correlation between the accuracy of forces and the accuracy of MD simulations. Recall that the energy and forces for all configurations in ANI-1ccx dataset have been evaluated at the DFT level of theory using the $\omega\text{B}97x$ functional 34, and we use this DFT-ANI-1ccx dataset only in Fig. 6. The teacher models are trained to the DFT ground truth energies. The first generation student models are trained to the ground truth DFT energies and the ensemble averaged forces from the teacher. The second generation student models are trained to the DFT energy and the ensemble averaged forces from the first generation students. Only the ‘baseline’ model is trained to the DFT energy and forces. Figure 6 summarizes the energy and force errors for the DFT-COMP6 benchmark dataset. The student models have much lower force errors than the teachers but are higher than the ‘baseline’ model which has access to the true energy and forces. The errors for $E$ and $\Delta E$ are comparable in the student and baseline models.
@@
To summarize, we introduce the EKD framework for achieving higher MLIP accuracy when training to datasets that include energies but not forces. An ensemble of teacher MLIPs are trained on the QC energies and then used to generate forces for all the configurations in the dataset. Then, student HIPNNs are trained to the ground truth QC energies and the ensemble-averaged forces from the teachers. The students exhibit up to $40\%$ improvements in the out-of-sample CC-COMP6 benchmarks for the energy and conformer energy differences, as well as more stable MD simulations. To probe the accuracy of the students’ forces, we apply the EKD to a dataset, where the energies and forces are computed at the DFT level of theory. The DFT forces are used only for testing, not for training, in the students, control and teachers. The student HIPNNs have lower errors with respect to the DFT forces when compared to the teachers. Additionally, the EKD workflow is effective even as the capacity gap between the teacher and student models grow. Although our workflow has the added cost of training an ensemble of teachers, it does not require any new expensive QC calculations beyond the original dataset needed to train the MLIPs, nor any exhaustive hyper-parameter tuning. Furthermore, our EKD workflow will be beneficial for reactive chemistry, where high fidelity QC methods are needed 47, 48. This is because the transition pathways generated with low fidelity methods, such as DFT, can show systematic deviations from high fidelity methods such as CCSDT 49, 50, 51, 52 or CASPT 53, 48 due to excessive charge delocalization 47.
@@
On a broader scope, our results serve as a proof of concept of model-agnostic knowledge distillation for regression tasks using deep neural networks. Previous KD methods based on intermediate outputs have shown limited success 19, 20 for regression, and feature matching KD 54, 19 approaches are dependent on the architectures of both the teachers and student models. This work paves the path towards model-agnostic KD methods, which will be relevant in constructing fast and accurate machine learning models.

### v3

# Ensemble Knowledge Distillation for Machine Learning Interatomic Potentials

@@
Sakib Matin sakibmatin@gmail.com Theoretical Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Emily Shinkle Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Yulia Pimonova Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Galen T. Craven Theoretical Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Aleksandra Pachalieva Earth and Environmental Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Ying Wai Li Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA    Kipton Barros [    Nicholas Lubbers Computer, Computational, and Statistical Sciences Division, Los Alamos National Laboratory, Los Alamos, NM 87545, USA
@@
The quality of machine learning interatomic potentials (MLIPs) strongly depends on the quantity of training data as well as the quantum chemistry (QC) level of theory used. Datasets generated with high-fidelity QC methods are typically restricted to small molecules and may be missing energy gradients, which make it difficult to train accurate MLIPs. We present an ensemble knowledge distillation (EKD) method to improve MLIP accuracy when trained to energy-only datasets. First, multiple teacher models are trained to QC energies and then generate atomic forces for all configurations in the dataset. Next, the student MLIP is trained to both QC energies and to ensemble-averaged forces generated by the teacher models. We apply this workflow on the ANI-1ccx dataset where the configuration energies computed at the coupled cluster level of theory. The resulting student MLIPs achieve new state-of-the-art accuracy on the COMP6 benchmark and show improved stability for molecular dynamics simulations.
@@
Machine learning models are a promising way to accelerate scientific simulations 1, 2, 3, 4. Machine learning interatomic potentials (MLIPs) 1, 2, 3 can emulate the potential energy surface of atomistic systems at dramatically reduced costs compared to reference quantum chemistry (QC) algorithms. MLIPs 5, 6, 2, 1, 3, 7, 8 are typically trained on QC datasets to map from atomic coordinates and species to configuration energy. The atomic forces can be predicted by the trained MLIPs, using automatic differentiation. There has been tremendous progress in the field of MLIP development, especially in terms of designing more expressive architectures, including equivariant descriptors 9, 10, 11 and transformer-based architectures 12. Furthermore recent works have focused on generating larger training data sets 13, 14, which often utilize different active learning protocols 15, 16, 17, 18 ,as well as scalable training 19. On the other hand, an under-explored area is the design of better training protocols, especially for low data regimes. Transfer learning 20 and multi-fidelity learning 21, 22 can be effective for small high fidelity datasets but they may require large amounts of data at a different level of theory to be successful.
Generating the data used to train MLIPs is far more computationally expensive than performing the training itself 18. It is especially computationally difficult to generate training data using high fidelity QC methods, such as Coupled Cluster, Configuration-Interaction, Quantum Monte Carlo, etc, because computational cost typically explodes with the number of electrons 23. Because of the much higher costs relative to density functional theory (DFT) calculations, datasets of high-fidelity QC are typically small, both in the size and number of molecular configurations. Furthermore, many high-fidelity QC codes provide only the total energy, but no gradients, which further hinders training of MLIPs 24. Due to the great expense of obtaining the gold standard of accuracy in quantum chemistry data, methods to make the most of limited datasets are of great utility.
@@
Recently, Knowledge Distillation (KD) for MLIPs 25, 26 has been shown to be an effective training protocol for existing datasets, without expensive pre-training. In the prototypical KD workflow 27, a single teacher model generates auxiliary outputs that augment the training of a student model in order to enhance speed 28, 25, memory requirements 29, and accuracy 30. References 31, 30, 32 have shown that multiple teachers can train a single student model to improve performance in classification tasks. KD has been applied to MLIPs to accelerate molecular dynamics (MD) simulations, by using intermediate outputs (e.g., atomic energies) 26, learned features 25, and hessians of the energy 33. Similar teacher-student training has also been applied to materials structure for property prediction tasks 34 and physics-constrained data augmentation 35. In related works (see Refs. 36, 37, 38), the teacher MLIP (trained on QC ground truth data) is used to generate synthetic data by running MD under different conditions. The student MLIPs is first trained to the synthetic data 36, 37, and optionally fine-tuned on other QC ground truth data 38. Existing studies have mostly focus on utilizing a single teacher and student workflow, even though using ensemble-averaged MLIPs in MD simulations has been shown to improve accuracy and stability 16. Only Ref. 39 studied knowledge distillation with an ensemble five graph neural networks (GNNs) trained on energy, forces and stress, which were evaluated at DFT level of theory. A GNN was chosen at random to generate new trajectories, for which the ensemble was used to predict energy, forces and stress, and then the model was fine-tuned to the mean predictions 39.
@@
Distinct from previous publications 25, 34, 39, 26, in our Ensemble Knowledge Distillation (EKD) for MLIPs workflow, a set of teacher MLIPs are trained on high fidelity data that only contains molecular energies. The trained teacher models can generate the atomic forces for all configurations in the dataset. The students are trained to the ground truth QC energies and the ensemble averaged forces from the teachers. Our workflow is outlined in Fig 1. We validate our EKD workflow on the ANI-1ccx dataset 20, 40 to show student models are more accurate and robust in our molecular dynamics tests than direct training. Our workflow establishes a new state-of-the-art accuracy for the ANI-1ccx dataset.
@@
Figure 1: Ensemble Knowledge Distillation for Hierarchically Interacting Particle Neural Network (HIPNN) 41, 11. The $k$ teacher models are trained on the reference quantum chemistry calculation of the energy and can generate forces (negative of the gradient of energy with respect to position). These forces are averaged over the ensemble of teachers and augment the student HIPNN training to improve the accuracy and robustness.
@@
In this letter, we apply the EKD method primarily to the Hierarchically Interacting Particle Neural Network (HIPNN) architecture 41, although this method can be readily applied to most existing MLIPs. HIPNN is a message-passing graph neural network 7, that can map atomic configurations to energy 41, forces 11, and to various chemical properties 42, 43, 44. HIPNN uses one-hot encoding, based on the atomic number, to initially featurize the local atomic environments. Then the interaction layers allow for mixing of atomic environments between neighbors (within a local cutoff) via message passing to refine the initial features 41. Using multiple interaction layers, $n_{\mathrm{Int}}>1$, implicitly accounts for some long-range physics 41, 11, 7. Although HIPNN 41 originally utilized scalar pair-wise distances between neighbors, the subsequent generalization—HIPNN with tensor sensitivity 11, 45—utilizes higher order Cartesian tensor products of the displacement vectors between neighboring atoms to construct more informative many-body descriptors. The hyper-parameter $l_{\mathrm{max}}$ corresponds to the highest order tensor used in the model. For $\ell_{\mathrm{max}}>0$, the model predictions are sensitive to the angles between neighboring atoms. The $\ell_{\mathrm{max}}=0$ HIPNN model coincides with the model developed in original publication 41. The atom layers (multi-layer perceptrons) predict the hierarchical contributions to the atomic energy, $\mathcal{\epsilon}_{i}$ which are then summed up to obtain the configuration, $E$. Automatic differentiation can be used to compute the forces on each atom $\boldsymbol{F}_{i}=-\boldsymbol{\nabla}_{i}E$. The hyper-parameters for the HIPNN models are given in Supporting Information.  3.1.
We validate our EKD workflow on the
@@
ANI-1ccx 20, 40 dataset, which consists of small organic molecules. The approximately $4.9\times 10^{5}$ molecular configurations in this dataset span $\mathrm{C,H,N,O}$ elements. The configurations have been down selected from the larger ANI-1x dataset (about 5 million configurations) using active learning 20, which utilized ensemble disagreement as the uncertainty metric. The non-equilibrium geometries are generated via normal mode sampling and short MD trajectories 15. The ANI-1ccx datasets include additional dihedral sampling of small molecules that is not present in the ANI-1x dataset. The dataset is available for download from Ref. 40. The configuration energies are computed at the coupled cluster with singles doubles, perturbative-triples, and complete basis set extrapolation [CCSD(T)/CBS] level of theory 46, 20 using the ORCA software 47. The dataset has also been computed at the DFT level of theory using the $\omega\text{B}97x$ functional, which is the same level of theory as the ANI-1x dataset 15.
@@
The COMP6 15, 20 is a challenging out-of-sample test for models trained to the ANI-1ccx and ANI-1x datasets. The configurations in the COMP6 are larger than the ANI-1x and ANI-1ccx training datasets, and provide a challenging extensibility test. For the ‘torsion’ 48 and ‘GDB 10-13’ data subsets, the energy and conformer energy differences have been computed at the CCSD(T)/CBS level of theory and we refer to these as the CC-COMP6 dataset. The conformational energy $\Delta E$ is the energy difference between all conformers for a given molecule in the benchmark. The conformers with energies at least $100\mathrm{kcal/mol}$ greater than the ground state are excluded, similar to the analysis in Ref. 20.
@@
Figure 2: Student HIPNNs have lower root-mean-squared-errors (RMSE) for energy $E$ and conformer energy differences $\Delta E$ compared to the teacher models. The error bars correspond to the standard deviation measured across $8$ models which differ by random weight initializations and data splits. The out-of-sample test sets “GDB 10-13” and “Torsion” 48 in panels A and B are subsets of the CC-COMP6 dataset 40.
@@
which contains the atomic positions $\boldsymbol{R}_{i}$ and species $Z_{i}$ for each configuration and the corresponding energy $E$. The choice of $8$ teachers was motivated by ensemble knowledge distillation for image classification 49, which highlighted diminishing returns beyond $10$ neural networks. In the Supplementary Information 1, we show that using a single teacher model is less effective than using an ensemble of $8$. All $8$ teacher models have the same architecture, but are initialized with different random weights and different data splits. We train the HIPNN models using stochastic gradient descent. The loss function consists of both error and regularization terms. The error loss $\mathcal{L}_{\mathrm{err}}$ is the sum of the root-mean-squared error (RMSE) and mean-absolute error (MAE) losses. The regularization term consists of the $L_{2}$ norm of model weights $\mathcal{L}_{L_{2}}$, which is commonly added to loss functions to reduce over-fitting, and the hierarchicality term $\mathcal{L}_{\mathrm{R}}$ 41, which is specific to HIPNN.
@@
Although the teacher models are trained only on molecular energies, they can predict the forces using automatic differentiation. The ensemble-averaged teacher forces for each atom,
@@
We use a loss scheduler where the $w_{\mathrm{F}}$ is dynamically updated during training. The value of $w_{\mathrm{F}}$ is larger during the early stages and slowly decreases to a smaller value. The student model training benefits from the local information provided by the $\overline{\boldsymbol{F}}_{i}$ in the early stages of training, followed by refinement on the QC configuration energies at the final stages. Similar loss scheduler 50 strategies have shown improved accuracy when training MLIPs. The weights of the loss function and scheduler are listed in the Supplementary Information 3.2. The force predictions from an ensemble of the first generation student models can be used to train the second generation of student HIPNNs.
We apply the EKD workflow to HIPNNs trained on the ANI-1ccx data set. The student HIPNNs achieve lower root-mean-squared error (RMSE) for energy $E$ and conformer energy differences $\Delta E$ in the out-of-sample CC-COMP6 benchmark compared to the set of teacher models in Fig. 2. The average error of the student HIPNNs is lower than the error of the ensembled predictions of teacher models. The ensembled teachers are more accurate than single teachers, but they have slower MD speed and increased memory requirements. By using EKD, we get a single student that does just as well or even better than the ensembled teachers. Thus, we can get all the gains in accuracy from the ensembled teacher model without the computational disadvantages of ensembled models.
@@
Figure 3: MD simulations driven by student HIPNNs are more stable than teacher models. The student and teachers are the same models as in Fig. 2. The MD simulations are performed with individual models, not ensembled predictions. Every data point is averaged over $8$ HIPNN models, where each MLIP is used to perform $10$ independent MD runs for a total of $80$ trajectories.
@@
To investigate how our EKD workflow is affected by the capacity gap between the teacher and student MLIPs, we vary the student models’ number of interaction layers $n_{\mathrm{Int}}$ and tensor sensitivity order $\ell_{\mathrm{max}}$. Increasing $n_{\mathrm{Int}}$ and $l_{\mathrm{max}}$ means that descriptors are more sensitive to many-body angular information about neighboring atoms, i.e., the resulting models have greater capacity 11. We fix the the width of the neural network, depth of the atom layers, and number of sensitivity functions, as well as the training hyper-parameters (number of epochs, learning rates, and optimizer) for all models considered. The control models have the same $n_{\mathrm{Int}}$ and $\ell_{\mathrm{max}}$ as the students but they are trained only on the QC energies. We set $n_{\mathrm{Int}}=2,\ell_{\mathrm{max}}=2$ for the teacher HIPNNs. Figure 4 shows that $E$ and $\Delta E$ RMSE of the student HIPNNs are consistently 10-30% lower than the control models on the CC-COMP6 benchmark.
@@
The accuracy of the forces of the student models in our EKD workflow is an important metric to analyze because of the strong correlation between the accuracy of forces and the accuracy of MD simulations. Recall that the energy and forces for all configurations in ANI-1ccx dataset have been evaluated at the DFT level of theory using the $\omega\text{B}97x$ functional 40, and we use this DFT-ANI-1ccx dataset only in Fig. 5. The teacher models are trained to the DFT ground truth energies. The first generation student models are trained to the ground truth DFT energies and the ensemble averaged forces from the teacher. The second generation student models are trained to the DFT energy and the ensemble averaged forces from the first generation students. Only the ‘baseline’ model is trained to the DFT energy and forces. Figure 5 summarizes the energy and force errors for the DFT-COMP6 benchmark dataset. The student models have much lower force errors than the teachers but are higher than the ‘baseline’ model which has access to the true energy and forces. The errors for $E$ and $\Delta E$ are comparable in the student and baseline models.
@@
To summarize, we introduce the EKD framework for achieving higher MLIP accuracy when training to datasets that include energies but not forces. An ensemble of teacher MLIPs are trained on the QC energies and then used to generate forces for all the configurations in the dataset. Then, student HIPNNs are trained to the ground truth QC energies and the ensemble-averaged forces from the teachers. The students exhibit up to $40\%$ improvements in the out-of-sample CC-COMP6 benchmarks for the energy and conformer energy differences, as well as more stable MD simulations. To probe the accuracy of the students’ forces, we apply the EKD to a dataset, where the energies and forces are computed at the DFT level of theory. The DFT forces are used only for testing, not for training, in the students, control and teachers. The student HIPNNs have lower errors with respect to the DFT forces when compared to the teachers. Additionally, the EKD workflow is effective even as the capacity gap between the teacher and student models grow. Although our workflow has the added cost of training an ensemble of teachers, it does not require any new expensive QC calculations beyond the original dataset needed to train the MLIPs, nor any exhaustive hyper-parameter tuning. Furthermore, our EKD workflow will be beneficial for reactive chemistry, where high fidelity QC methods are needed 52, 53. This is because the transition pathways generated with low fidelity methods, such as DFT, can show systematic deviations from high fidelity methods such as CCSDT 54, 55, 56, 57 or CASPT 58, 53 due to excessive charge delocalization 52.
@@
On a broader scope, our results are an important example of model-agnostic knowledge distillation for regression tasks using deep neural networks 33. Previous KD methods based on intermediate outputs have shown limited success 25, 26 for regression, and feature matching KD 59, 25 approaches are dependent on the architectures of both the teachers and student models. This work paves the path towards model-agnostic KD methods, which will be relevant in constructing fast and accurate machine learning models.
@@
1 Knowledge Distillation using a single teacher
We perform knowledge distillation with a single teacher model and a single student model for the Ani-1ccx dataset. The student model is trained to the ground truth energies and the forces from the teacher model. For the teacher model, the “GDB 10-13” $E$ and $\Delta E$ RMSE are $3.31\mathrm{kcal/mol}$, and $2.45\mathrm{kcal/mol}$ respectively. The student HIPNN performs marginally better, and the ‘GDB 10-13” $E$ and $\Delta E$ RMSE are $2.93\mathrm{kcal/mol}$, and $2.11\mathrm{kcal/mol}$ respectively. We find that using only a single teacher model is less beneficial than using an ensemble, as seen in Fig. 2.
2 EKD across different MLIP architectures
We implement the EKD method with different MLIP architectures for teachers and students. We use TorchANI-1ccx MLIPs from Refs. 60 and 61, which were first pre-trained to the nearly 5 million configurations of the DFT ANI-1x dataset, and fine-tuned to the coupled cluster ANI-1ccx dataset 20. We compare the $\Delta E$ RMSE of the “GDB 10-13” subset of the CC-COMP6 benchmark in Fig. 6. We find that the using a teacher torchANI model does benefit a student HIPNN model, compared to direct training (teacher HIPNNs). Ultimately, we find that using HIPNN MLIPs for both teacher and student gives lowest error, which we attribute to the fact the teacher HIPNNs have lower errors than the teacher TorchANI models.